Find the present value of $10,000 received at the start of every year for 20 years if the interest rate is J1 = 12% p.a. and if the first payment of $10,000 is received at the end of 10 years
1
Expert's answer
2020-06-14T17:21:12-0400

The present value will be calculated as follows:

The interest for the 10 years:

"\text{Simple interest}=100000×10×12100=120000"


Therefore, the amount will be:

"\text{Amount} = 10000+12000=22000"




After 20yrs another payment received value of principal is 10000*20 = 200000

"\text{Simple interest}=20000×20×12100=48,000"


"\text{Amount} = 48,000+20,000=68,000"

Present value will therefore be:

"\text{Present value}=22000\times \dfrac{1}{(1+0.12)^{10}}+68000\times \dfrac{1}{(1+0.12)^{20}}"

"\text{Present value}=14,132.75"